OpenStudy (layshay12):

Which number line best shows the position of square root of 3?

OpenStudy (layshay12):

Number line with numbers negative 4, negative 3, negative 2, negative 1, 0, 1, 2, 3, 4 labeled. A dot is made at 1.7 and labeled as square root of 3. Number line with numbers negative 4, negative 3, negative 2, negative 1, 0, 1, 2, 3, 4 labeled. A dot is made at 2.8 and labeled as square root of 3. Number line with numbers negative 4, negative 3, negative 2, negative 1, 0, 1, 2, 3, 4 labeled. A dot is made at 3 and labeled as square root of 3. Number line with numbers negative 4, negative 3, negative 2, negative 1, 0, 1, 2, 3, 4 labeled. A dot is made at 2.2 and labeled as square root of 3.

OpenStudy (layshay12):

Please help?? Will medal and fan

OpenStudy (rayaz):

\[\sqrt{3}<\sqrt{4}\ thus\ \sqrt{3}<2\] \[\sqrt{3}>\sqrt{2}>\sqrt{1}\ \therefore \sqrt{3}>\sqrt{2}>1\ thus \sqrt{3}>1\] From this we can conclude that \[1<\sqrt{3}<2\] The number line that satisfies the inequality is the answer. (the first one)

OpenStudy (layshay12):

Thanks that was right!!! Can you help me with another question

OpenStudy (rayaz):

sure

OpenStudy (layshay12):

What is the simplified expression for 4 to the power of negative 3 multiplied by 3 to the power of 4 multiplied by 4 to the power of 2 whole over 3 to the power of 5 multiplied by 4 to the power of negative 2 ? 3 over 4 4 over 3 4 to the power of 2 over 3 3 to the power of 3 over 4

OpenStudy (sshayer):



OpenStudy (layshay12):

Thx. How about question 2

OpenStudy (rayaz):

Gather all like terms \[\frac{ 4^{-3}\times4^{2} }{ 4^{-2} }\times \frac{ 3^{4} }{ 3^{5} }\] Apply the laws of exponents i.e \[x^a\times\ x^b = x^{a+b}\]\[\frac{ x^a }{ x^b } = x ^{a-b}\]

OpenStudy (layshay12):

Okay thx. You have been really helpful. I'm sure to get a good grade